Fermi National Laboratory
 

Accelerator Update for May 22, 2006 - May 23, 2006

 

Monday May 22

The day shift began with Tevatron separator testing.

Technicians and experts halted separator tests and accessed the TeV sectors A-E and the Transfer Hall. They came out of access near the end of the day shift.

Permission to run beam was given to the 8 GeV line, Main Injector (MI), and Recycler.

Permission was given to run the Pelletron.

Permission was given to MiniBooNE to take beam.

MI was successfully ramping by 3 PM.

Operations established beam to Booster and MI at 4:36 PM. Experts began tuning and conducting studies. MI had a problem capturing and accelerating beam so experts conducted 8 GeV studies.

TeV separator conditioning continued.

Operators discovered that there was a soft ground fault in the TeV somewhere between sectors B2 and C2.

Tuesday May 23

The midnight shift began with MI/8 GeV studies and TeV separator conditioning.

The 8 GeV line studies finished.

A Recycler expert began setting up their machine.

A Linac RF station (LRF1) started having problems. An expert came in to investigate.

Operations prepared the TeV for access at 7:25 AM.

The Plans for Today

The plans are to allow TeV accesses to complete work and to give priority to the MI high-field orbit measurements.

Machine Reports

Linac

No report.

Booster

No report.

MI

Techs repaired the coasting beam valve. Experts managed to get beam around MI on the first pulse, but beam fell out of the machine as they increased its energy. Experts think there might be a problem with MECAR. Experts made MI-8 lattice measurements. They found three bad MI BPMs and were having problems with a Multiwire. They want to conduct high-field measurements.

Pbar

The AP1 line needs a vacuum leak check. The A30 bake was completed and the tank might be cool enough to power later this morning. The Debuncher is cold and experts will start extraction tests. The enclosures have been cleaned. CUB suffered a LCW pump failure – Pbar has no backup pump at this time. If there is beam available, they will be ready for it by Wednesday night.

TeV

All the separators are on and conditioning with very few sparks. There are some controls hardware problems with the separators. The Tevatron Electron Lens (TEL1) installation should be completed today.

Recycler

Experts are waiting for beam. Techs are working on the electron cooling system and hope to be ready for electrons by the end of the week.

Switchyard 120

Experts are conducting walkthroughs today. They are concerned about locating the P3 line ground fault quickly.

MiniBooNE

Experts want beam to their target today to test equipment.

NuMI

Experts have started up equipment and will run for four days. They still need their enclosures cleaned. Electrician work continues. They will be ready for beam by Friday.

Collider Detector at Fermilab

Experts ran cosmic ray tests and found noise problems.

D-Zero

They are working on their detector today and on making configuration changes tomorrow. They will have their detector closed up by Thursday.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/28/2007    email Fermilab